X is a random variable taking values 0 and 1 with respective probabilities q and p. A sample X1, X2, ..., Xn of size n is taken from the distribution. If r = summation i = 1 to n(Xi) , show that (r + 1))/(n + 1) is a biased estimator of p. Also show that the bias of the estimator tends to zero
as n tends to infinity.
1
Expert's answer
2020-09-13T18:15:36-0400

"\frac{r+1}{n+1}" is an estimator of the parameter "p". We will show that "\frac{r+1}{n+1}" is a biased estimator of "p" using the definiton of a biased estimator.

"E(\frac{r+1}{n+1})=\frac{1}{n+1}E(r+1)=\frac{1}{n+1}E(\sum_{i=1}^n X_i+1)=\\
=(E(X_i)=p)=\frac{1}{n+1}(np+1)\neq p."

So "\frac{r+1}{n+1}" is a biased estimator of "p".

"\text{Bias}(\frac{r+1}{n+1})=E(\frac{r+1}{n+1})-p=\frac{np+1}{n+1}-p=\frac{1-p}{n+1}.\\
\frac{1-p}{n+1}\rightarrow 0\text{ as } n\rightarrow \infty."